how do you find a vertical and horizontal intercept? (-6,6) (5,2)
would I be right to say
6-2 = 4
-6-5 + -11

Answer by stanbon(75887)

how do you find a vertical and horizontal intercept? (-6,6) (5,2)
would I be right to say
6-2 = 4
-6-5 + -11
That would help you find the slope of the line through the
two points.
m=-4/11
Now find "b" in y=mx+b by substituting 2=-(4/11)(5)+b
22/11=-20/11 + b
b=42/11
Your equation is y=(-4/11)x+42/11
Multiply thru by 11 to get
11y=-4x+42
To find the x-intercept let y=0:
You get x=42/4=21/2
So the x-intercept is (21/2,0)
To find the y-intercept let x=0:
You get y=42/11
So the y-intercept is (0,42/11)
Cheers,
Stan H.